OpenStudy (angelab97):

At central high school, 40% of the members of the Spanish club are boys. How many members does the club have if 30 of the members are girls?

Directrix (directrix):

If 40% of the members are boys, then 60% of the members are girls. Note that 40% + 60% = 100% which is the entire club. If T represents the total membership of the club, then 60% of T is equal to 30. .6T = 30 Solve for T @AngelaB97 Post what you get, and we'll compare answers. Okay?

OpenStudy (angelab97):

i got 20 because if 0.6/30=0.40(x)

OpenStudy (angelab97):

@Directrix

Directrix (directrix):

.6T = 30 T = 30 divided by .6 T = 30 divided by 6/10 T = 30 times 10/6 T = 300/6 T = ?

OpenStudy (angelab97):

is there an easier way. I'm having trouble understanding

Directrix (directrix):

Divide 30 by .6 --------------- I did not follow this -->0.6/30=0.40(x) I do know that 20 is not correct for the members in the club. We know the club has 30 girls so how could the total membership be only 20?

OpenStudy (angelab97):

OMG thanks. I found that there's 20 boys and that added to 30 equals 50 total members.

Directrix (directrix):

Oh, okay, you found the boys. 50 total is what I got, too.